At a school, 52 % of people are taking the advanced functions class, 36 % of people are taking the calculus class, and 21 % of people are taking neither. If a student is chosen at random, what is the probability that they are in calculus, but not advanced functions?
Answer #1

P(advanced functions class) = 0.52

P(calculus class) = 0.36

P(neither class) = 0.21

P(advanced functions class or calculus class) = 1 - P(neither class) = 1 - 0.21 = 0.79

P(advanced functions class or calculus class) = P(advanced functions class) + P(calculus class) - P(advanced functions class and calculus class)

or, 0.79 = 0.52 + 0.36 - P(advanced functions class and calculus class)

or, P(advanced functions class and calculus class) = 0.09

P(calculus, but not advanced functions) = P(calculus class) - P(advanced functions class and calculus class)

                                                               = 0.36 - 0.09

                                                               = 0.27